# Crystal-field magnetostriction of the spin ice under ultrahigh magnetic fields

### v1

@@ III.1 High-field magnetization and magnetostriction
Figure 2: Magnetic-field dependence of (a) magnetization $M$ and (b) longitudinal magnetostriction $\Delta L/L\parallel[111]$ in Ho2Ti2O7 for $B\parallel[111]$ in ultrahigh magnetic fields. Black symbols in (a) show the $M$$B$ data measured at $T=4.2$ K using the MPMS, purple symbols show the data measured at $T_{\rm ini}=4.2$ K using the nondestructive pulsed magnet, and green symbols show the data measured at $T_{\rm ini}=5$ K using the HSTC system. The insets of panels (a) and (b) show the field derivatives of magnetization $dM/dB$ and magnetostriction $d(\Delta L/L)/dB$, respectively, for the HSTC data. The critical field of the CF level crossing $B_{\rm cf}$ is defined using the field-down sweep data.
Figure 2(b) shows the longitudinal magnetostriction curve for $B\parallel[111]$ measured up to 120 T at $T_{\rm ini}=5$ K using the HSTC system, along with the curve measured up to 55 T at $T_{\rm ini}=4.2$ K using the nondestructive pulsed magnet. Initially, $\Delta L/L$ increases and reaches a maximum around 30 T. Subsequently, $\Delta L/L$ reverses and dramatically decreases from 30 to 80 T, where the field derivative of the magnetostriction $d(\Delta L/L)/dB$ exhibits a peak at $B_{\rm cf}\sim 65$ T for the field-down sweep [see the inset of Fig. 2(b)]. Finally, $\Delta L/L$ is nearly saturated at $\sim$80 T. The observed change in sample length across $B_{\rm cf}$, amounting to $\Delta L/L\sim 5\times 10^{-4}$, highlights the strong CF effect on the crystal structure of Ho2Ti2O7.
@@ III.2 Magnetocaloric effect (MCE)
Comparing the three $M$$B$ curves shown in Fig. 2(a), it is evident that the initial rise in magnetization at low magnetic fields becomes more gradual as the field sweep becomes faster. This trend could be attributed to two factors: sample heating and slow spin relaxation. We will discuss the spin dynamics in Sec. III.3. To verify the sample heating, we performed MCE measurements for $B\parallel[111]$ up to 55 T in the nondestructive pulsed magnet. The field dependence of the sample temperature $T(B)$, measured at various initial temperatures ($T_{\rm ini}$’s) under quasi-adiabatic conditions, are shown by the blue curves in Fig. 3. The $T(B)$ curves for field-up and field-down sweeps overlap for all the measured $T_{\rm ini}$’s. The absence of hysteresis ensures that nearly adiabatic conditions were achieved. These $T(B)$ curves can hence be regarded as isentropic, meaning that the total entropy, comprising both lattice entropy and magnetic entropy, remains constant.
For $T_{\rm ini}=5$ K, $T(B)$ approaches 30 K at 30 T, indicating a significant reduction in magnetic entropy Kittaka_2018 . This occurs because the highly degenerate “2-in–2-out” states are destroyed by a magnetic field as weak as 2 T, stabilizing the nondegenerate “3(1)-in–1(3)-out” state. At higher initial temperatures, the increase in $T(B)$ is less pronounced due to greater thermal fluctuations and larger lattice heat capacity. After $T(B)$ reaches its maximum around 35 T, the sample gradually cools as the magnetic field increases up to 55 T. This cooling behavior indicates an increase in magnetic entropy, suggesting the onset of CF level hybridization.
We also measured the sample temperature changes at $T_{\rm ini}=4.2$ K in a liquid-4He environment, i.e., under nonadiabatic conditions, as shown by the red curve in Fig. 3. Surprisingly, $T(B)$ reaches 28 K during the field-up sweep, which is only slightly lower than in the quasi-adiabatic case. These MCE results indicate that sample heating effects are unavoidable when measuring Ho2Ti2O7 under pulsed high magnetic fields. We infer that, in the magnetization and magnetostriction data presented in Fig. 2, the sample temperature is likely between 20 and 30 K at fields above 10 T.
Figure 3: Magnetocaloric effect in Ho2Ti2O7 for $B\parallel[111]$ measured at various $T_{\rm ini}$’s under quasi-adiabatic conditions (blue), and at $T_{\rm ini}=4.2$ K under nonadiabatic conditions (red).
@@ III.3 Spin dynamics
Our high-field data also provide valuable insights into the dynamical properties of the magnetic state. As for the experimental data obtained in the millisecond-duration pulsed magnetic fields [Figs. 7(a) and 7(b)], both the $M$$B$ curve up to 62 T and the $\Delta L/L$$B$ curve up to 55 T exhibit small hysteresis (Fig. 2). These hysteresis are likely due to the lower sample temperature during the field-down sweep compared to the field-up sweep, as suggested by our nonadiabatic MCE data (Fig. 3). We note that the hysteresis loop in the $\Delta L/L$$B$ curve cannot be attributed to slow spin dynamics because the sign of the hysteresis is opposite to the delayed response. This suggests that the timescale of the spin dynamics is faster than the millisecond range.
@@
In contrast, the $M$$B$ curve up to 130 T, obtained in the microsecond-duration pulsed magnetic fields [Fig. 7(c)], exhibits significant hysteresis [Fig. 2(a)]. Given that no hysteresis is observed in the quasi-adiabatic MCE (Fig. 3), the observed hysteresis can be attributed to slow spin relaxation, which originates from nonequilibrium processes of monopole formation or annihilation TStoter_2020 . Notably, the hysteresis observed in the $\Delta L/L$$B$ curve up to 120 T is much less pronounced compared to that in the $M$$B$ curve (Fig. 2). This suggests that the relaxation time associated with CF striction, which is dominant in the high-field magnetostriction, is likely faster than the microsecond range. This timescale is consistent with previous studies based on AC susceptibility, neutron spin echo, and $\mu$SR experiments GEhlers_2003 ; BTomasello_2019 .
@@ III.5 Numerical calculations
Figure 5: Calculation results of (a) $M$$B$ curves and (b) $\Delta L/L$$B$ curves for the Hamiltonian in Eq. (5), obtained by mean-field approach. The inset of panel (a) shows an enlarged view of the $M$$B$ curve at 5 K in the low-field regime, clarifying a metamagnetic behavior from the “2-in–2-out” to the “3(1)-in–1(3)-out” state. In the inset of panel (b), the circle and square symbols represent the contributions from exchange striction ($\Delta L_{\rm ex}/L$) and crystal field (CF) striction ($\Delta L_{\rm cf}/L$), respectively, to the total magnetostriction at 5 K.
Having Confirmed that our simulations accurately describe the magnetization of Ho2Ti2O7 across the entire field range, we now turn to the longitudinal magnetostriction. According to Eq. (7), the total magnetostriction is composed of the CF striction ($\Delta L_{\rm cf}/L$) and exchange striction ($\Delta L_{\rm ex}/L$). The field dependence of $\Delta L_{\rm cf}/L$ and $\Delta L_{\rm ex}/L$ at 5 K is plotted in the inset of Fig. 5(b). The combination of a positive change in $\Delta L_{\rm ex}/L$ in the low-field regime and a negative change in $\Delta L_{\rm cf}/L$ in the high-field regime results in a hump structure in the total magnetostriction curve. Notably, the rounded shape of the experimentally observed hump in the $\Delta L/L$$B$ curve at $T_{\rm ini}=5$ K [Fig. 2(b)] is presumably due to sample heating, which is consistent with the calculated $\Delta L/L$$B$ curve at 20 K. Our calculations predict that the hump in $\Delta L/L$ shifts to a lower magnetic field with increasing temperature and eventually disappears above 60 K. This behavior also agrees well with the experimental observations, as shown in Fig. 4(a).
Figure 6: Magnetic-field dependence of (a) magnetization $M$ and (b) longitudinal magnetostriction $\Delta L/L\parallel[111]$ in Pr2Zr2O7 for $B\parallel[111]$ in ultrahigh magnetic fields. Black symbols in (a) show the $M$$B$ data measured at $T=4.2$ K using the MPMS, purple symbols show the data measured at $T_{\rm ini}=4.2$ K using the nondestructive pulsed magnet, and green symbols show the data measured at $T_{\rm ini}=5$ K using the HSTC system. The inset of panel (b) show the transverse magnetostriction data measured at $T_{\rm ini}=4.2$ K using the nondestructive pulsed magnet.
@@ IV Conclusion
We have thoroughly investigated the static and dynamic magnetoelastic properties in the spin-ice compound Ho2Ti2O7 through magnetostriction measurements under ultrahigh magnetic fields. Our observations reveal a large anisotropic magnetostriction occurring at the CF level crossing around $B_{\rm cf}\sim 65$ T, where the longitudinal magnetostriction amounts to $\Delta L/L\sim-5\times 10^{-4}$. This is an order of magnitude larger than the exchange striction observed in the low-field regime. Complementary MCE measurements reveal a dramatic sample-heating effect exceeding 20 K in pulsed high-field experiments, and also highlight the onset of CF level hybridization above 30 T, which would correlate with the turnaround behavior of the magnetostriction curve. By varying the magnetic field sweep rates, we distinguish between the fast dynamics of the CF-phonon coupling and the slow dynamics of spin-spin correlations.
To model the CF striction, we developed a comprehensive Hamiltonian based on a point-charge model that includes CF-phonon coupling. Our numerical calculations based on the mean-field approach successfully account for the experimental magnetostriction data across the entire field range. This study paves the way for understanding the intricate role of the CF-phonon coupling in the crystal structure of a broad range of $4f$-based magnets.
@@ ACKNOWLEDGMENTS
This work was partly supported by the JSPS KAKENHI Grants-In-Aid for Scientific Research (No. 20J10988 and No. 24H01633), JST-ASPIRE Program (JPMJAP2317), and JST-MIRAI Program (JPMJMI20A1). N.T. was supported by the Alexander von Homboldt foundation. M.G. was supported by the JSPS through a Grant-in-Aid for JSPS Fellows. We appreciate the stimulating discussions with R. Moessner, R. Klingeler, M. Gingras, T. Fennell, and M. Udagawa.
@@ Appendix B Phonon and CF-phonon coupling terms
Here, $\bm{p}_{i}$ and $m_{i}$ denote the momentum and mass, respectively, of the atom at position $i$. The displacement vector of the atom at position $i$ is represented by $\bm{U}_{i}$. The difference between the undisplaced lattice positions is given by $\bm{R}_{ij}\equiv\bm{R}_{j}-\bm{R}_{i}$, where $\bm{R}_{j}$ ($\bm{R}_{i}$) represents the position of atom at site $j$ ($i$). $k_{ij}$ is the longitudinal atomic spring constants, which is determined based on the Born-von Karman model, assuming simple distance-dependent elastic constant (see Appendix C for details). Note that the sum over indices $i$ and $j$ counts each spring twice, thus a factor of $1/2$ is included before the sum. The elastic constants are then used to define the phonon-strain coupling constant $G_{\rm mix}$ as follows:
@@
where $a_{0}$ is the Bohr radius, and the indices $\alpha,\beta,\gamma=x,y,z$ refer to an Euclidean coordinate system parallel to the crystal axes, i.e., $x\parallel a$, $y\parallel b$ and $z\parallel c$.
@@
Importantly, $G{\rm{}_{cfph}}$ affects the magnitude of $\Delta L/L$. We scaled $G{\rm{}_{cfph}}$ obtained from McPhase by a factor of 0.16 to match the calculated $\Delta L/L$ with the experimental one. Although the elastic constants $c^{\alpha\beta}$ also affect the magnitude of $\Delta L/L$, their values were fixed based on prior experimental data, as detailed in Appendix C.
@@
where $A$ and $D$ are set to 0.4 and 3.49, respectively, resulting in $J_{\rm ex}=0.1$ K and $dJ_{\rm ex}/dr=1.03$ $\AA$ at the nearest neighbour distance $R_{\rm nn}=3.54$$\AA$ Baroudi_2015 . It is important to constrain $dJ_{\rm ex}/dr$ to a positive value for Ho2Ti2O7 in order to explain the positive longitudinal magnetostriction in the low-field regime.
Appendix C Elastic constant

### v3

@@ III.1 High-field magnetization and magnetostriction
Figure 2: Magnetic-field dependence of (a) magnetization $M$ and (b) longitudinal magnetostriction $\Delta L/L\parallel[111]$ in Ho2Ti2O7 for $B\parallel[111]$ in ultrahigh magnetic fields. Black symbols in (a) show the $M$$B$ data measured at $T=4.2$ K using the MPMS, purple symbols show the data measured at $T_{\rm ini}=4.2$ K using the nondestructive pulsed magnet, and green symbols show the data measured at $T_{\rm ini}=5$ K using the HSTC system. The insets of (a) and (b) show the field derivatives of magnetization $dM/dB$ and magnetostriction $d(\Delta L/L)/dB$, respectively, for the HSTC data. The critical field of the CF level crossing $B_{\rm cf}$ is defined using the field-down sweep data.
Figure 2(b) shows the longitudinal magnetostriction curve for $B\parallel[111]$ measured up to 120 T at $T_{\rm ini}=5$ K using the HSTC system, along with the curve measured up to 55 T at $T_{\rm ini}=4.2$ K using the nondestructive pulsed magnet. Initially, $\Delta L/L$ increases and reaches a maximum around 30 T. Subsequently, $\Delta L/L$ reverses and dramatically decreases from 30 to 80 T, where the field derivative of the magnetostriction $d(\Delta L/L)/dB$ exhibits a peak at $B_{\rm cf}\sim 65$ T for the field-down sweep [see the inset of Fig. 2(b)]. Finally, $\Delta L/L$ is nearly saturated at $\sim$80 T. The observed change in sample-length across $B_{\rm cf}$, amounting to $\Delta L/L\sim 5\times 10^{-4}$, highlights the strong CF effect on the crystal structure of Ho2Ti2O7.
@@ III.2 Magnetocaloric effect (MCE)
Comparing the three $M$$B$ curves shown in Fig. 2(a), it is evident that the initial rise in magnetization at low magnetic fields becomes more gradual as the field sweep becomes faster. This trend could be attributed to two factors: sample heating and slow spin relaxation. We will discuss the spin dynamics in Sec. III.3. To verify the sample heating, we performed MCE measurements for $B\parallel[111]$ up to 55 T in the nondestructive pulsed magnet. The field dependence of the sample temperature $T(B)$, measured at various initial temperatures ($T_{\rm ini}$’s) under quasiadiabatic conditions, are shown by the blue curves in Fig. 3. The $T(B)$ curves for field-up and field-down sweeps overlap for all the measured $T_{\rm ini}$’s. The absence of hysteresis ensures that nearly adiabatic conditions were achieved. These $T(B)$ curves can hence be regarded as isentropic, meaning that the total entropy, comprising both lattice entropy and magnetic entropy, remains constant.
For $T_{\rm ini}=5$ K, $T(B)$ approaches 30 K at 30 T, indicating a significant reduction in magnetic entropy Kittaka_2018 . This occurs because the highly degenerate 2-in–2-out states are destroyed by a magnetic field as weak as 2 T, stabilizing the nondegenerate 3(1)-in–1(3)-out state. At higher initial temperatures, the increase in $T(B)$ is less pronounced due to greater thermal fluctuations and larger lattice heat capacity. After $T(B)$ reaches its maximum around 35 T, the sample gradually cools as the magnetic field increases up to 55 T. This cooling behavior indicates an increase in magnetic entropy, suggesting the onset of CF level hybridization.
We also measured the sample temperature changes at $T_{\rm ini}=4.2$ K in a liquid-4He environment, i.e., under nonadiabatic conditions, as shown by the red curve in Fig. 3. Surprisingly, $T(B)$ reaches 28 K during the field-up sweep, which is only slightly lower than in the quasiadiabatic case. These MCE results indicate that sample heating effects are unavoidable when measuring Ho2Ti2O7 in pulsed high magnetic fields. We infer that, in the magnetization and magnetostriction data presented in Fig. 2, the sample temperature is likely between 20 and 30 K at fields above 10 T.
Figure 3: Magnetocaloric effect in Ho2Ti2O7 for $B\parallel[111]$ measured at various $T_{\rm ini}$’s under quasiadiabatic conditions (blue), and at $T_{\rm ini}=4.2$ K under nonadiabatic conditions (red).
@@ III.3 Spin dynamics
Our high-field data also provide valuable insights into the dynamic properties of the magnetic state. In the experimental data obtained using the millisecond-duration pulsed magnetic fields [Figs. 7(a) and 7(b)], both the $M$$B$ curve up to 62 T and the $\Delta L/L$$B$ curve up to 55 T exhibit small hysteresis (Fig. 2). The hysteresis is likely caused by a lower sample temperature during the field-down sweep compared to the field-up sweep, as suggested by our nonadiabatic MCE data (Fig. 3). We note that the hysteresis loop in the $\Delta L/L$$B$ curve cannot be attributed to slow spin dynamics because the sign of the hysteresis is opposite to what would be expected from a delayed response. Therefore, the small hysteresis observed in millisecond-duration pulsed magnetic-fields suggests the timescale of spin relaxation is faster than the millisecond range, which will be discussed in more detail in the next paragraph.
@@
In contrast, the $M$$B$ curve up to 130 T, obtained in the microsecond-duration pulsed magnetic fields [Fig. 7(c)] under quasi-adibatic condition, exhibits significant hysteresis [Fig. 2(a)]. Given that no hysteresis is observed in the quasiadiabatic MCE (Fig. 3), the observed hysteresis can be attributed to slow spin relaxation, which mainly originates from nonequilibrium processes of monopole formation or annihilation TStoter_2020 . Together with the magnetization data from millisecond-pulsed fields, the spin dynamics appears to be faster than the millisecond but slower than or comparable to the microsecond regime, consistent with previous studies based on neutron spin echo and ac susceptibility experiments GEhlers_2003 . However, inelastic neutron scattering experiments have reported a different timescale of 10 ns, suggesting the presence of faster processes JPClancy_2009 . The reason for such discrepancy has been addressed in Ref. Quilliam_2011 .
Notably, the hysteresis observed in the $\Delta L/L$$B$ curve up to 120 T is much less pronounced compared to that in the $M$$B$ curve (Fig. 2), a difference that cannot be explained by temperature effects, as both experiments were performed under quasiadiabatic condition. The overall sample-length change of Ho2Ti2O7 in high magnetic fields primarily originates from CF-striction, with a smaller contribution from exchange striction (see Sec. III.5), making CF-striction the main source of hysteresis. This suggests that the dynamics of CF-striction is likely within or faster than the microsecond range. It is important to note that potential contributions from slower spin-lattice dynamics cannot be excluded, given the absence, to the best of our knowledge, of previous research on such interactions in Ho2Ti2O7, as these dynamics should occur beyond the timescale of our pulsed-field experiments. Further investigation into the dynamics of spin-lattice interactions in Ho2Ti2O7 would be valuable for future work.
@@ III.5 Numerical calculations
Figure 5: Calculation results of (a) $M$$B$ curves and (b) $\Delta L/L$$B$ curves for the Hamiltonian in Eq. (5), obtained by mean-field approach. The inset of (a) shows an enlarged view of the $M$$B$ curve at 5 K in the low-field regime, clarifying a metamagnetic behavior from the “2-in–2-out” to the “3(1)-in–1(3)-out” state. In the inset of (b), the circle and square symbols represent the contributions from exchange striction ($\Delta L_{\rm ex}/L$) and crystal-field (CF) striction ($\Delta L_{\rm cf}/L$), respectively, to the total magnetostriction at 5 K.
Having Confirmed that our simulations accurately describe the magnetization of Ho2Ti2O7 across the entire field range, we now turn to the longitudinal magnetostriction. According to Eq. (7), the total magnetostriction is composed of the CF striction ($\Delta L_{\rm cf}/L$) and exchange striction ($\Delta L_{\rm ex}/L$). The field dependence of $\Delta L_{\rm cf}/L$ and $\Delta L_{\rm ex}/L$ at 5 K is plotted in the inset of Fig. 5(b), where $\Delta L_{\rm cf}/L$ becomes more prominent in the high-field regime. The combination of a positive change in $\Delta L_{\rm ex}/L$ in the low-field regime and a negative change in $\Delta L_{\rm cf}/L$ in the high-field regime results in a hump structure centering around 30 T in the total magnetostriction curve. Notably, the rounded shape of the experimentally observed hump in the $\Delta L/L$$B$ curve at $T_{\rm ini}=5$ K [Fig. 2(b)] is presumably due to sample heating, which is consistent with the calculated $\Delta L/L$$B$ curve at 20 K. Our calculations predict that the hump in $\Delta L/L$ shifts to a lower magnetic field with increasing temperature and eventually disappears above 60 K. This behavior also agrees well with the experimental observations, as shown in Fig. 4(a).
Figure 6: Magnetic-field dependence of (a) magnetization $M$ and (b) longitudinal magnetostriction $\Delta L/L\parallel[111]$ in Pr2Zr2O7 for $B\parallel[111]$ in ultrahigh magnetic fields. Black symbols in (a) show the $M$$B$ data measured at $T=4.2$ K using the MPMS, purple symbols show the data measured at $T_{\rm ini}=4.2$ K using the nondestructive pulsed magnet, and green symbols show the data measured at $T_{\rm ini}=5$ K using the HSTC system. The inset of (b) shows the transverse magnetostriction data measured at $T_{\rm ini}=4.2$ K using the nondestructive pulsed magnet.
@@ IV Conclusion
We have thoroughly investigated the static and dynamic magnetoelastic properties in the spin-ice compound Ho2Ti2O7 through magnetostriction measurements under ultrahigh magnetic fields. Our observations reveal a large anisotropic magnetostriction occurring at the CF level crossing around $B_{\rm cf}\sim 65$ T, where the longitudinal magnetostriction amounts to $\Delta L/L\sim-5\times 10^{-4}$. Complementary MCE measurements reveal a dramatic sample-heating effect exceeding 20 K in pulsed high-field experiments, and also highlight the onset of CF level hybridization above 30 T, which would correlate with the turnaround behavior of the magnetostriction curve. By varying the magnetic field sweep rates, we distinguish between the fast dynamics of the CF-phonon coupling and the slower dynamics of spin-spin correlations, associated with monopole creation and annihilation.
To model the CF striction, we developed a comprehensive Hamiltonian based on a point-charge model that includes CF-phonon coupling. Our numerical calculations based on a mean-field approach successfully reproduce the experimental magnetostriction data across the entire field range, showing a dominant contribution from CF striction $\Delta L/L_{\rm cf}$ and a much smaller contribution from exchange striction $\Delta L/L_{\rm ex}$. This study paves the way for understanding the intricate role of the CF-phonon coupling in the crystal structure of a broad range of $4f$-based magnets.
@@ ACKNOWLEDGMENTS
This work was partly supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via TRR 360 (project No. 492547816), JSPS KAKENHI Grants-In-Aid for Scientific Research (Grants No. 20J10988 and No. 24H01633), JST-ASPIRE Program (JPMJAP2317), and JST-MIRAI Program (JPMJMI20A1). N.T. was supported by the Alexander von Humboldt foundation. M.G. was supported by the JSPS through a Grant-in-Aid for JSPS Fellows. We gratefully acknowledge the enduring support of T. Sakakibara and the stimulating discussions with R. Moessner, R. Klingeler, M. Gingras, T. Fennell, and M. Udagawa.
@@ Appendix B Phonon and CF-phonon coupling terms
Here, $\bm{p}_{i}$ and $m_{i}$ denote the momentum and mass, respectively, of the atom at position $i$. The displacement vector of the atom at position $i$ is represented by $\bm{U}_{i}$. The difference between the undisplaced lattice positions is given by $\bm{R}_{ij}\equiv\bm{R}_{j}-\bm{R}_{i}$, where $\bm{R}_{j}$ ($\bm{R}_{i}$) represents the position of atom at site $j$ ($i$). $k_{ij}$ is the longitudinal atomic spring constants, which is determined based on the Born-von Karman model, assuming simple distance-dependence (see Appendix C for details). The spring constants are then used to define the phonon-strain coupling constant $G_{\rm mix}$ as follows:
@@
where $a_{0}$ is the Bohr radius, and the indices $\alpha,\beta,\gamma=x,y,z$ refer to an Euclidean coordinate system parallel to the crystal axes, i.e., $x\parallel a$, $y\parallel b$ and $z\parallel c$. To be noted, in Eq. (7), $\alpha$ and $\beta$ are used in Voigt notation ($\alpha,\beta=1,...,6$), and are defined differently than in this context.
@@
Again, in Eq. (7), $\alpha$ and $\beta$ are used in Voigt notation ($\alpha,\beta=1,...,6$), and are defined differently than in this context. Importantly, $G{\rm{}_{cfph}}$ affects the magnitude of $\Delta L/L$. We scaled $G{\rm{}_{cfph}}$ obtained from McPhase by a factor of 0.16 to match the calculated $\Delta L/L$ with the experimental one. Although the elastic constants $c^{\alpha\beta}$ also affect the magnitude of $\Delta L/L$, their values were fixed based on prior experimental data, as detailed in Appendix C.
@@
where $A$ and $D$ are set to 0.4 and 3.49, respectively, resulting in $J_{\rm ex}=0.1$ K and $dJ_{\rm ex}/dr=1.03$ $\AA$ at the nearest-neighbor distance $R_{\rm nn}=3.54$$\AA$ Baroudi_2015 . It is important to constrain $dJ_{\rm ex}/dr$ to a positive value for Ho2Ti2O7 in order to explain the positive longitudinal magnetostriction in the low-field regime.
Appendix C Elastic constants